OpenStudy (anonymous):

A homeowner wants several small dormers built on his roof to admit light. They will be in the shape of isosceles triangles whose sides will exceed their base by 5 inches. If the perimeter is 100 inches, what are the lengths of the three sides? The base is___, and the equal sides are____.

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

oh so simple

OpenStudy (mayankdevnani):

in isosceles triangle two sides are equal

OpenStudy (anonymous):

a+B+C

OpenStudy (anonymous):

i got that i did it and for some reason my answers were wrong

OpenStudy (mayankdevnani):

so x+x+5=100 2x=95 x=95/2=47.5

OpenStudy (mayankdevnani):

so three sides are 47.5,47.5 and 5 inches

OpenStudy (anonymous):

that's not the answer

OpenStudy (mayankdevnani):

@Claire4christ its right answer

OpenStudy (anonymous):

the answer is not 47

OpenStudy (anonymous):

two equal sides, and both exceed the base by 5 base is x side is x + 5 perimeter is x + (x+5) + (x +5) = 100 3x +10 = 100 3x = 90 x = 30 base = 30, sides = 35

OpenStudy (anonymous):

i got the same and 35 wasnt right

OpenStudy (mayankdevnani):

oh!! i got it

OpenStudy (anonymous):

@JakeV8 you are right

OpenStudy (anonymous):

so, base = 30, sides = 35 is right? When the problem says the sides exceed the base by 5, that is a little vague and open to different interpretations.

OpenStudy (anonymous):

glad to help... :)

OpenStudy (anonymous):

your just forgat to put inches